# Quantization-Aware Neuromorphic Architecture for Efficient Skin Disease Classification on Resource-Constrained Devices

### v2

@@ IV-E Inference Speed and Energy Consumption
IV-F Ablation Study of Model Components
Table IV summarizes a stepwise ablation in which we start from a compact baseline and incrementally enable Ghost, ECA, SE, quantization-aware head with the same augmentation policy, embedding-space SMOTE, and on-chip incremental learning. All experiments use the same input resolution, optimizer, schedules, and 8-bit quantization configuration to isolate the effect of each component. Metrics are computed on the same test split.
The drop observed for “Baseline + Ghost” is expected under a strictly width-preserving replacement of standard convolutions by Ghost modules. Ghost generates a fraction of channels through inexpensive linear transforms of a reduced set of primary channels. When used alone and early in the stack, this reduces the number of learned base filters that directly capture fine dermatoscopic structures such as pigment networks, streaks, and globules. The cheap transforms introduce correlated feature maps that do not add discriminative power without a mechanism to suppress redundancy. Under 8-bit activation clipping, these correlated responses compress the effective dynamic range and make the batch-norm folding more sensitive to local contrast, which in turn degrades minority-class recall. The combination of early downsampling and reduced primary channels further limits high-frequency detail preserved from $64\times 64$ inputs. As a result, the test accuracy decreases relative to the baseline when Ghost is introduced in isolation.
Adding ECA recovers and surpasses the baseline by providing lightweight channel recalibration that suppresses redundant ghosted maps and reweights informative channels. The subsequent SE block contributes a complementary global channel prior that stabilizes inter-class boundaries and improves calibration. Enabling the quantization-aware head with the same augmentation policy improves the match between integer ranges at the readout and the upstream bounded activations, which tightens the loss landscape after BN folding. Injecting SMOTE in the embedding space raises coverage of minority manifolds without perturbing low-level image statistics and yields consistent gains in macro metrics. Finally, on-chip incremental learning adapts the last-stage parameters to post-conversion activation statistics and to the deployment data distribution, producing the highest accuracy and F1. We also observed that widening the primary branch of Ghost or delaying its use to later stages can mitigate the initial drop, but we kept the compact configuration to satisfy on-chip memory and latency constraints and relied on attention and calibration to recover accuracy.
IV-G Classification Performance on Clinical Dataset
@@
We evaluated the proposed neuromorphic skin lesion classification system and several representative CNN architectures, all converted to SNNs and deployed on the Akida hardware platform. Table V presents the Top-1 accuracy and macro F1 score for each model under the same deployment conditions. Our model achieves the highest accuracy and macro F1 score among all tested methods, confirming its effectiveness and robustness for neuromorphic inference in clinical scenarios.
V Conclusion
We presented QANA, a quantization-aware neuromorphic pipeline for skin lesion classification on edge hardware. Under a fixed $64\times 64$ input and identical preprocessing across methods, QANA reaches 91.6% Top-1 and 82.4% macro F1 on HAM10000 and 90.8%/81.7% on a clinical dataset. The compiled SNN runs at 1.5 ms on chip and 1.7 mJ per image on Akida, measured as power above idle. These results indicate that conversion-preserving design and deployment-aware training enable accurate and efficient on-device dermatology inference.
VI Acknowledgement
This research was supported by the National Natural Science Foundation of China (Nos. 62172336 and 62032018). The authors gratefully acknowledge BrainChip Holdings Ltd. for providing technical support and the Akida AKD1000 hardware platform, whose powerful neuromorphic computing capabilities enabled strong performance of the SNN model. The authors also extend their appreciation to Dr. Atif Mansoor and his teams for their preliminary contributions to this research.

### v3

@@ 4.4 Classification Performance on HAM10000
4.5 Inference Speed and Energy Consumption
We quantitatively evaluated inference latency and energy consumption against conventional CNN baselines and other converted SNN architectures. All SNN models were deployed on the Akida AKD1000 PCIe board, while CNN baselines were tested on the NVIDIA RTX 3090 GPU and the Intel Core i9-12900K CPU. Latency is reported for batch size $1$ with a fixed input pipeline, after a warm-up phase, and averaged over 10,000 images. Energy per image is computed as the mean device power during inference multiplied by the average per-image latency, using on-board power monitoring for Akida, NVIDIA-smi for GPU, and Intel RAPL for CPU. For each model, the reported inference latency corresponds to the average per-image processing time over 10,000 test samples. Energy consumption per image was measured as the mean device power during inference multiplied by the average inference time, using on-board power monitoring for Akida and NVIDIA-smi for GPU and Intel RAPL for CPU.
To ensure a fair and hardware-consistent comparison, all tested CNN architectures were converted to SNNs using our pipeline before deployment on Akida. Table 3 presents a comprehensive summary of inference latency and energy consumption across all evaluated models and platforms, as well as the relative reduction of these metrics for SNNs on Akida compared to their CNN GPU and CPU implementations. QANA achieves the lowest inference latency and energy consumption among the evaluated architectures. When deployed as an SNN on Akida, it processes an image in 1.5 ms and consumes 1.7 mJ, which corresponds to a 94.6% latency reduction and a 99.0% energy reduction relative to its CNN implementation on the RTX 3090.
4.6 Ablation Study of Model Components
Table 4 reports the ablation results on the HAM10000 test set, illustrating the contribution of each core module to the overall performance. Modules were incrementally enabled to measure their isolated and cumulative effects on classification metrics, including accuracy, recall, precision, F1 score, and AUC-ROC (%). The results demonstrate that each module provides consistent improvements, with the complete model achieving the highest accuracy and F1 score.
4.7 Classification Performance on Clinical Dataset
@@
We evaluated the proposed neuromorphic skin lesion classification system and several representative CNN architectures, all converted to SNNs and deployed on the Akida hardware platform. Table 5 presents the Top-1 accuracy and macro F1 score for each model under the same deployment conditions. Our model achieves the highest accuracy and macro F1 score among all tested methods, confirming its effectiveness and robustness for neuromorphic inference in clinical scenarios.
5 Conclusion
This paper presented QANA, a quantization-aware neuromorphic framework for on-device skin lesion classification. Experiments on HAM10000 and a real-world clinical dataset show that QANA maintains competitive diagnostic accuracy while enabling efficient SNN deployment on BrainChip Akida. On Akida hardware, the converted model achieves 1.5 ms per-image latency and 1.7 mJ per image, supporting real-time inference under strict energy and privacy constraints.